OpenStudy (anonymous):

Evaluate negative one thousand three hundred thirty one raised to the two thirds

OpenStudy (anonymous):

-1335^(2/3) = ...

OpenStudy (campbell_st):

well you need to evaluate \[\sqrt[3]{1331^2}\]

OpenStudy (anonymous):

(-1331)^2/3

OpenStudy (anonymous):

correct

OpenStudy (campbell_st):

or an easy way is \[1331 = 11^3 \] so your problem is \[(11^3)^\frac{2}{3} \] multiply the powers for the solution.

OpenStudy (campbell_st):

well make your answer negative...

OpenStudy (anonymous):

-123 is what I got @campbell_st

OpenStudy (campbell_st):

lol... not close so the problem is \[-1331^\frac{2}{3} = (-11^3)^\frac{2}{3}\] the index law for power of a power is \[(x^a)^b = x^{a \times b}\] so you have \[-11^{3 \times \frac{2}{3}} = \] you just need the answer

OpenStudy (anonymous):

Lol so confusing, 121 ? @campbell_st

OpenStudy (campbell_st):

thats it